OpenStudy (anonymous):

Use L'Hopitals Rule to Calculate: Lim (tan(2x))^x X ->O^+

OpenStudy (anonymous):

so far i have

OpenStudy (amistre64):

i spose thats a good start

OpenStudy (amistre64):

tan = sin/cos right?

OpenStudy (amistre64):

not sure if that helps with the ^x; might have to do some ln to it

OpenStudy (amistre64):

y = (tan(2x))^x ln(y) = x ln(tan(2x)) y'/y = x' ln(tan(2x)) + x ln'(tan(2x)) y'/y = ln(tan(2x)) + x/tan(2x) *tan(2x)' y'/y = ln(tan(2x)) + 2x sec^2(2x)/tan(2x) not really sure how that would help tho lol

OpenStudy (amistre64):

1 --- cs 2x sec(x)csc(x) ... just playing

OpenStudy (anonymous):

\[\lim_{x\to 0}x\ln(\tan(2x))\] is in the form \(0\times -\infty\) standard trick is to write it as a fraction as \[\frac{\ln(\tan(2x))}{\frac{1}{x}}\] and use l'hopital

OpenStudy (anonymous):

first time you get \[\frac{-2x^2\cos(2x)}{\sin(2x)}\] which is still \(\frac{0}{0}\) so repeat the process. second time you get 0

OpenStudy (anonymous):

therefore your limit is \[e^0=1\]

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

yw i left the algebra for you

OpenStudy (anonymous):

ok cool :)